# Forgetting is Everywhere

### v1

@@ Appendix C Thought experiments
We present a series of thought experiments designed to stress-test potential definitions of forgetting in general learning systems. Each scenario illustrates a conceptual edge case that any robust definition of forgetting should naturally handle. For each scenario, we provide a consistency verdict, indicating whether our framework classifies the learner to be forgetting or not, and explain the intuition behind this judgement.
@@
Consistency verdict: no forgetting. Since the state remains constant across any number of updates, the induced futures distributions are identical. A learner that never changes its state can never forget.
@@
Consistency verdict: 0-bit does not forget, the 5-bit stack does. The predictive distribution of the stack is always the returned bit of the address being accessed at each step. A 0-bit stack’s state never changes, so its future remains constant. The 5-bit stack, however, overwrites past bits when it is full; its predictive distribution changes as bits are dropped, hence it forgets. Forgetting only exists in systems capable of learning.
@@
Consistency verdict: no forgetting. The predictive distributions of such a learner are a Dirac at the retrieved value (or null if unassigned), and the induced futures remain identical over time. Thus, the learner never forgets.
@@
Consistency verdict: no forgetting. Under regular environmental updates, the clock advances its state (its notion of “time”) by one unit per tick. However, under consistency updates, we propagate the learner’s state forward using its own current beliefs rather than the actual environmental progression. Since the clock believes time remains fixed at its last known state, it will continue to predict the exact time indefinitely. Consequently, its futures remain invariant under consistency updates. The clock does not forget.
@@
Consistency verdict: no forgetting (on ignored steps). When no update occurs, the state remains identical, so induced futures distributions are unchanged. Forgetting can only occur during state-altering updates.
@@
Consistency verdict: no forgetting. In expectation, the induced futures after any number of consistent updates remain invariant. The randomness of selection does not imply loss of information about possible futures.
@@
Consistency verdict: no forgetting. Similar to the clock, the consistency update believes time remains fixed at its last known state; thus, the learner does not flip its predictions, and its state remains fixed. Consequently, its futures remain invariant under consistency updates, and the flipper does not forget.
@@
Consistency verdict: no forgetting. Similar to the clock, the consistency update believes time remains fixed at its last known state; thus, the learner does not flip its predictions during consistent updates, and its state and predictive distributions remain fixed. Consequently, its futures remain invariant under consistency updates, and the learner does not forget.
@@
Consistency verdict: this forgets. Even though the test data were unseen, the learner’s induced futures change when its performance on those data changes. Thus, the learner has forgotten: forgetting occurs whenever the previously supported predictive capabilities vanish, regardless of whether the corresponding data have been observed.
@@
Consistency verdict: no forgetting. The next $k$ observations may either be a 0 or a $1$, drawn from any distribution. After $k$ updates, the learner’s predictive distribution is deterministically either a 0 or a $1$. Although the state continues to evolve across the induced futures, the expected predictive distribution remains the same across them. Therefore, no forgetting occurs.
@@
Consistency verdict: no forgetting. Under consistent updates, future updates are performed relative to the learner’s current beliefs rather than the actual environment. This distinction is what prevents surprise from being mistaken for forgetting. Since the learner’s predictive distribution already encodes its own uncertainty over future outcomes, an unexpected tail does not constitute loss of information; it is a valid observation under the learner’s model. Consequently, the expected futures distribution before and after the $k$ updates remains identical, and no forgetting is detected.
@@
Consistency verdict: no forgetting. As in C.11, future updates are computed relative to the learner’s current predictive distribution rather than the actual environment. Because updates reflect the learner’s own beliefs, encountering a new minimum does not imply a loss of prior information. In expectation, the futures distributions before and after $k$ updates remain identical, indicating that the learner has not forgotten.
@@ Appendix D Empirical considerations
In this section, we detail the approach used to estimate the forgetfulness measure $\Gamma_{k}(t)$. $\Gamma_{k}(t)$ captures the extent to which the learner’s induced futures at time $t$ remain consistent with those obtained after $k$ simulated updates. While this quantity cannot be computed exactly, it can be approximated using a particle-based Monte Carlo scheme. At a given time, we clone the learner into $M$ replicas and propagate each replica forward under $k$ steps of simulated interaction with its own induced futures. This yields a set of post-update models $\{Z_{t+k-1}^{(m)}\}_{m=1}^{M}$, whose mixture distribution provides an approximation to the consistency condition:
@@
We compare the learner’s distribution on a reference trajectory of simulated futures with this mixture distribution. The divergence between the two provides an empirical estimate of $\Gamma_{k}(t)$.
@@
Figure 6 compares a single-layer neural network’s futures distribution at the current training step with the same model but after $40$ additional updates on futures data. This allows us to visualise the divergence between the two induced futures distributions. This difference quantifies the model’s propensity to forget, as defined in Definition 4.2.
Figure 6: Comparison of initial and 40-step induced futures. Axes show the input space around the two-moon binary classification task. Example datapoints are overlaid and coloured by target class ($\text{red}=0$, $\text{blue}=1$). The background grid is shaded according to the classifier’s logit values, from red (0) to blue (1), with white indicating uncertainty at 0.5. Left: The initial induced futures distribution $q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ serves as the baseline for consistent updates. Right: The $k=40$ induced futures distribution $q^{*}_{k}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$, showing how the predictive model is likely to evolve over 40 steps. The divergence in Definition 4.2 is computed between these two distributions.
@@ Appendix E Additional results
E.1 Unforgetful learners
Figure 7: Self-consistent learners do not forget. Axes showing observations $(x,y)$ are shaded white, and parameter axes $(w_{0},w_{1})$ are shaded grey. Top row: The same four observations are presented to a linear regression learner in different orders. Second row: A full Bayesian posterior represents a complete summary of all observations and satisfies the $k$-step consistency condition, implying that no forgetting has occurred. The permutation invariance of the learner in i.i.d. settings is a corollary of this property, due to Bayesian commutative updates. Bottom rows: Two constrained learners – a Gaussian variational posterior with diagonal covariance and a point estimate updated by gradient descent – violate self-consistency. Their updates alter induced futures; therefore, previously supported capabilities may be lost. The final column shows the resulting posterior predictive distributions.
We illustrate forgetting behaviour using Bayesian linear regression, where the exact Bayesian learner is analytically tractable, self-consistent, and therefore does not forget.
At time $t$, the learners beliefs are captured by its predictive distribution $p(X_{t+1:\infty}\mid\theta_{t},X_{0:t})$. Let $X_{1:t}$ denote the observations to time $t$, and let $\theta$ denote the parameters. The Bayesian posterior after observing $X_{1:t}$ is
$p(\theta\mid X_{1:t})\propto p(\theta)\prod_{i=1}^{n}p(X_{i}\mid\theta).$ (21)
Since multiplication commutes, the posterior is invariant to the order of observations:
$p(\theta\mid X_{1:t})=p(\theta\mid X_{\pi(1)},\dots,X_{\pi(t)})\quad\forall\text{ permutations }\pi.$ (22)
Therefore, exact Bayesian updates are permutation-invariant in i.i.d. settings.
Self-consistency of Bayesian updates.
Exact Bayesian learning forms a projective family. Marginalising the posterior after a hypothetical future observation recovers the current posterior:
$p(\theta\mid X_{1:t})=\int p(\theta\mid X_{1:t+1})p(X_{t+1}\mid X_{1:t})\,\mathrm{d}X_{t+1}.$ (23)
Equivalently, conditioning on a future observation and then marginalising over its prior predictive distribution returns the same belief as not conditioning at all. Thus, conditioning and marginalising commute. This is equivalent to the $k$-step self-consistency.
Self-consistency ensures that every predictive capability supported by $p(\theta\mid X_{1:t})$ remains supported after a Bayesian update unless contradicted by data. The learner preserves its prior predictive beliefs in expectation; therefore, the consistency condition is never violated, and no forgetting occurs.
Approximate procedures such as diagonal Gaussian variational inference or gradient-based point estimates do not satisfy the self-consistency condition. Their updates depend on the order of observations and yield posteriors that cannot be recovered by marginalising futures. As a result, previously supported predictions may be lost, indicating that forgetting has occurred. The predictive and parameter distributions for all learners are shown in Figure 7.
E.2 Impact of changes to algorithms
@@
Here we expand on the empirical results presented in §5, examining how the dynamics of forgetting vary with model architecture, hyperparameters, and optimisation settings.
@@
Figure 8: Impact of model size on forgetting dynamics. Plots illustrating the impact of varying numbers of hidden-layer parameters in a single-hidden-layer neural network on a regression task with 40 training datapoints (details in Appendix F). Left: Forgetfulness as a function of the number of updates $k$, showing how the learner’s propensity to forget evolves over update steps. Middle: Forgetfulness throughout training, highlighting changes in the learner’s propensity to forget over time. Right: Training loss curves, comparing learning dynamics across model sizes. We observe that forgetting dynamics are strongly influenced by model size: forgetting increases with model size until the learner’s parameter count approaches or exceeds the task’s effective size, at which point forgetting reduces again. These dynamics have a major impact on training efficiency, as shown in Figure 4.
@@
Figure 8 examines how model size affects forgetting and learning efficiency on the regression task detailed in Appendix F. Each plot compares networks with different hidden-layer sizes, trained on 40 datapoints that span both under- and over-parameterised tasks. We find that forgetting generally increases with model size, peaking near the point where the number of parameters matches the task’s effective complexity, and then decreases again in the highly overparameterised task. This inverted trend mirrors the behaviour of training efficiency discussed in Figure 4.
@@
In Figure 9, we vary the momentum coefficient in stochastic gradient descent while training on the same regression task. The propensity to forget exhibits periodic fluctuations whose amplitude increases with higher momentum values. Both the mean level of forgetting and the oscillatory behaviour grow with momentum, indicating a trade-off between stability and adaptivity. The most efficient training occurs around a momentum of 0.9, where moderate levels of forgetting coincide with stable and fast convergence.
@@
Figure 10 studies the influence of batch size on forgetting dynamics. Smaller minibatches lead to higher update variability, resulting in larger oscillations and a greater propensity to forget as training progresses. As batch size increases, forgetting decreases and training becomes more stable, plateauing at around one-quarter of the training dataset size. Beyond this, forgetting remains non-zero and is very similar across batch sizes greater than one-quarter of the training dataset.
@@
Finally, Figure 11 examines how model architecture influences forgetting on a high-dimensional classification task (CIFAR-10). We compare logistic regression, a multilayer perceptron (MLP), a convolutional neural network (CNN) and a ResNet. CNNs and ResNets exhibit considerably higher forgetting than logistic regression and MLPs, yet they achieve better task performance. This indicates that the most effective deep learning models do not necessarily minimise forgetting; instead, they maintain a balance where moderate forgetting supports continued adaptation and improved generalisation.
@@
Together, these results (Figure 8, Figure 9, and Figure 10) show that forgetting is a pervasive property of deep learning systems, shaped by both the update dynamics and the learner’s hyperparameters. While high levels of forgetting can destabilise training, moderate levels appear beneficial: enabling adaptability while preserving some past information. The relationship between forgetting and learning efficiency is thus nuanced: efficient learning requires not the absence of forgetting but its regulation.

### v2

@@ Appendix C Thought experiments
We present a series of thought experiments designed to stress-test potential definitions of forgetting in general learning systems. Each scenario illustrates a conceptual edge case that any robust definition of forgetting should naturally handle. For each scenario, we provide a self-consistency verdict, indicating whether our framework classifies the learner to be forgetting or not, and explain the intuition behind this judgement.
@@
Self-consistency verdict: no forgetting. Since the state remains constant across any number of updates, the induced futures distributions are identical. A learner that never changes its state can never forget.
@@
Self-Consistency verdict: 0-bit does not forget, the 5-bit stack does. The predictive distribution of the stack is always the returned bit of the address being accessed at each step. A 0-bit stack’s state never changes, so its future remains constant. The 5-bit stack, however, overwrites past bits when it is full; its predictive distribution changes as bits are dropped, hence it forgets. Forgetting only exists in systems capable of learning.
@@
Self-Consistency verdict: no forgetting. The predictive distributions of such a learner are a Dirac at the retrieved value (or null if unassigned), and the induced futures remain identical over time. Thus, the learner never forgets.
@@
Self-Consistency verdict: no forgetting. Under regular environmental updates, the clock advances its state (its notion of “time”) by one unit per tick. However, under self-consistent updates, we propagate the learner’s state forward using its own current beliefs rather than the actual environmental progression. Since the clock believes time remains fixed at its last known state, it will continue to predict the exact time indefinitely. Consequently, its futures remain invariant under self-consistent updates. The clock does not forget.
@@
Self-Consistency verdict: no forgetting (on ignored steps). When no update occurs, the state remains identical, so induced futures distributions are unchanged. Forgetting can only occur during state-altering updates.
@@
Self-Consistency verdict: no forgetting. In expectation, the induced futures after any number of consistent updates remain invariant. The randomness of selection does not imply loss of information about possible futures.
@@
Self-Consistency verdict: no forgetting. Similar to the clock, the self-consistent update believes time remains fixed at its last known state; thus, the learner does not flip its predictions, and its state remains fixed. Consequently, its futures remain invariant under self-consistent updates, and the flipper does not forget.
@@
Self-Consistency verdict: no forgetting. Similar to the clock, the self-consistent update believes time remains fixed at its last known state; thus, the learner does not flip its predictions during consistent updates, and its state and predictive distributions remain fixed. Consequently, its futures remain invariant under self-consistent updates, and the learner does not forget.
@@
Self-Consistency verdict: this forgets. Even though the test data were unseen, the learner’s predictive distribution changes when its performance on those data changes. Thus, the learner has forgotten: forgetting occurs whenever the previously supported predictive capabilities vanish, regardless of whether the corresponding data have been observed.
@@
Self-Consistency verdict: no forgetting. The next $k$ observations may either be a 0 or a $1$, drawn from any distribution. After $k$ updates, the learner’s predictive distribution is deterministically either a 0 or a $1$. Although the state continues to evolve across the induced futures, the expected predictive distribution remains the same across them. Therefore, no forgetting occurs.
@@
Self-Consistency verdict: no forgetting. Under consistent updates, future updates are performed relative to the learner’s current beliefs rather than the actual environment. This distinction is what prevents surprise from being mistaken for forgetting. Since the learner’s predictive distribution already encodes its own uncertainty over future outcomes, an unexpected tail does not constitute loss of information; it is a valid observation under the learner’s model. Consequently, the expected futures distribution before and after the $k$ updates remains identical, and no forgetting is detected.
@@
Self-Consistency verdict: no forgetting. As in § C, future updates are computed relative to the learner’s current predictive distribution rather than the actual environment. Because updates reflect the learner’s own beliefs, encountering a new minimum does not imply a loss of prior information. In expectation, the futures distributions before and after $k$ updates remain identical, indicating that the learner has not forgotten.
@@ Appendix D Empirical considerations
In this section, we detail the approach used to estimate the forgetfulness measure $\Gamma_{k}(t)$. $\Gamma_{k}(t)$ captures the extent to which the learner’s induced predictive distribution at time $t$ remains consistent with those obtained after $k$ simulated updates. While the simulated marginalisation $q_{k}^{*}(\ldots)$ cannot be computed exactly, it can be approximated using a particle-based Monte Carlo scheme. At a given time, we clone the learner into $M$ replicas and propagate each replica forward under $k$ steps of simulated interaction with its own induced futures. This yields a set of post-update models $\{Z_{t+k-1}^{(m)}\}_{m=1}^{M}$, whose mixture distribution provides an approximation to the self-consistency condition:
@@
We compare the learner’s distribution on a reference trajectory of simulated futures with this mixture distribution. The divergence between the two provides an empirical estimate of $\Gamma_{k}(t)$. This is detailed in Algorithm 1, provided below.
@@
Figure 6 compares a single-layer neural network’s futures distribution at the current training step with the same model but after $40$ additional updates on futures data. This allows us to visualise the divergence between the two induced futures distributions. This difference quantifies the model’s propensity to forget, as defined in § 4.2.
Figure 6: Comparison of initial and 40-step induced futures. Axes show the input space around the two-moon binary classification task. Example datapoints are overlaid and coloured by target class ($\text{red}=0$, $\text{blue}=1$). The background grid is shaded according to the classifier’s logit values, from red (0) to blue (1), with white indicating uncertainty at 0.5. Left: The initial induced futures distribution $q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ serves as the baseline for consistent updates. Right: The $k=40$ induced futures distribution $q^{*}_{k}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$, showing how the predictive model is likely to evolve over 40 steps. The divergence in § 4.2 is computed between these two distributions.
D.1 Supervised classification example
Consider a supervised classification setting with a neural network learner. To approximate future inputs $X_{t:\infty}$, we sample uniformly over the empirical distribution of inputs observed thus far. This is acquired by sampling $k$, $X$ values from a held-out validation set, $X_{t:t+k-1}$, and using the remaining inputs to evaluate predictive distributions, approximating $X_{t+k:\infty}$.
We first compute the predictive distribution of the current model, $Z_{t-1}$.
$q(Y_{t+k:\infty}\mid Z_{t-1},X_{t+k-1:\infty}),$ (56)
which will act as the reference distribution in the KL divergence.
To form the $k$-step simulated marginalisation, we perform a Monte Carlo estimate: make $N$ independent particle copies of the learner and, for each particle $n$ and for $s=1,\dots,k$, we do the following:
$\begin{aligned} &\text{(sample target)}&&Y_{t+s}^{(n)}\sim q_{f}(\cdot\mid Z_{t+s-1}^{(n)},X_{t+s-1}^{(n)}),\\ &\text{(sample input)}&&X_{t+s}^{(n)}\sim q_{e}(\cdot\mid H_{0:t+s-1},Y_{t+s}^{(n)}),\\ &\text{(update state)}&&Z_{t+s}^{(n)}\sim u(\cdot\mid Z_{t+s-1}^{(n)},X_{t+s}^{(n)},Y_{t+s}^{(n)}).\end{aligned}$ (57)
After $k$ steps, each particle yields a predictive distribution over the future inputs,
$q_{k}^{(n)}(Y_{t+k:\infty}\mid Z_{t+k-1}^{(n)},X_{t+k-1:\infty}^{(n)}).$ (58)
We form the uniform mixture to acquire the simulated marginalisation,
$q_{k}^{*}(Y_{t+k:\infty}\mid Z_{t+k-1},X_{t+k-1:\infty}).$ (59)
We can then compute the learner’s propensity to forget:
$\Gamma_{k}(t)=D_{\mathrm{KL}}\left(q(Y_{t+k:\infty}\mid Z_{t-1},X_{t+k-1:\infty})\,\|\,q_{k}^{*}(Y_{t+k:\infty}\mid Z_{t+k-1},X_{t+k-1:\infty})\right).$ (60)
This process is described in Algorithm 1.
Algorithm 1 Computing the propensity to forget of a learner in a supervised classification setting.
Learner with state $Z_{t-1}$, update function $u$, inputs $X_{t:\infty}$, # of particles NN, horizon $k$, history $H_{0:t-1}$
Propensity to Forget $\Gamma_{k}(t)$
Compute initial predictive distribution $q(Y_{t+k:\infty}\mid Z_{t-1},X_{t+k-1:\infty})$
Initialise NN particle copies of the learner: $\{Z_{t-1}^{(n)}\}_{n=1}^{N}$
for each particle $n=1$ to NN do
  for $s=1$ to $k$ do
   Sample target $Y_{t+s}^{(n)}\sim q_{f}(\cdot\mid Z_{t+s-1}^{(n)},X_{t+s-1}^{(n)})$
   Sample input $X_{t+s}^{(n)}\sim q_{e}(\cdot\mid H_{0:t+s-1},Y_{t+s}^{(n)})$
   Update particle state: $Z_{t+s}^{(n)}\sim u(\cdot\mid Z_{t+s-1}^{(n)},X_{t+s}^{(n)},Y_{t+s}^{(n)})$
  end for
  Compute predictive distribution for future inputs $q_{k}^{(n)}(Y_{t+k:\infty}\mid Z_{t+k-1}^{(n)},X_{t+k-1:\infty}^{(n)})$
end for
Form a mixture of the NN particle predictions $q_{k}^{*}(Y_{t+k:\infty}\mid Z_{t+k-1},X_{t+k-1:\infty})$
Compute $\Gamma_{k}(t)=\mathrm{D}_{\mathrm{KL}}\left(q(Y_{t+k:\infty}\mid Z_{t-1},X_{t+k-1:\infty})\,\|\,q_{k}^{*}(Y_{t+k:\infty}\mid Z_{t+k-1},X_{t+k-1:\infty})\right)$
return $\Gamma_{k}(t)$
Computational considerations.
The primary computational cost of this procedure stems from propagating $N$ independent particles over $k$ synthetic update steps, each involving a forward pass and parameter update of the learner. This can be significant for large neural networks or long consistency horizons $k$. Two sources of approximation error arise in practice. First, the predictive distributions $q_{f}$ and $q_{k}^{*}$ are evaluated only on a finite validation set, introducing sampling error that decreases with the number of held-out inputs. Second, the mixture $q_{k}^{*}$ is only a Monte Carlo approximation to the expectation over the learner’s self-generated future. Both sources can be reduced by increasing the size of the validation set or the number of particles, though at the cost of increased computational load.
@@ Appendix E Additional results
Figure 7: Optimal forgetting is not necessarily zero. We examine how a regression learner’s training efficiency and propensity to forget vary under different hyperparameter settings. Left: Increasing the batch size reduces forgetting while improving training efficiency. Forgetting plateaus once the batch size approaches the dataset size (40 datapoints), at which point training efficiency increases with minimal further reduction in forgetting. Right: Reducing the noise in the dataset leads to substantially higher training efficiency; however, in low-noise regimes, the learner becomes more sensitive to self-consistency perturbations and therefore forgets more despite performing better on the task.
E.1 Forgetting in supervised learning
@@
Here we expand on the empirical results presented in §5, examining how the dynamics of forgetting vary with model architecture, hyperparameters, and optimisation settings. In addition to the training efficiency analysis in Figure 4, we observe that hyperparameters can simultaneously shape both training efficiency and forgetting in Figure 7.
@@
Figure 8: Impact of model size on forgetting dynamics. Plots illustrating the impact of varying numbers of hidden-layer parameters in a single-hidden-layer neural network on a regression task with 40 training datapoints (details in § F). Left: Forgetfulness as a function of the number of updates $k$, showing how the learner’s propensity to forget evolves over update steps. Middle: Forgetfulness throughout training, highlighting changes in the learner’s propensity to forget over time. Right: Training loss curves, comparing learning dynamics across model sizes. We observe that forgetting dynamics are strongly influenced by model size: forgetting increases with model size until the learner’s parameter count approaches or exceeds the task’s effective size, at which point forgetting reduces again. These dynamics have a major impact on training efficiency, as shown in Figure 4.
@@
Figure 8 examines how model size affects forgetting and learning efficiency on the regression task detailed in § F. Each plot compares networks with different hidden-layer sizes, trained on 40 datapoints that span both under- and over-parameterised tasks. We find that forgetting generally increases with model size, peaking near the point where the number of parameters matches the task’s effective complexity, and then decreases again in the highly overparameterised task. This inverted trend mirrors the behaviour of training efficiency discussed in Figure 4.
@@
In Figure 9, we vary the momentum coefficient in stochastic gradient descent while training on a regression task. The propensity to forget exhibits periodic fluctuations, increasing with higher momentum values. Both mean forgetting, and the oscillatory behaviour grow with momentum, indicating a trade-off between stability and adaptivity. The most efficient training occurs around a momentum of 0.9, where moderate levels of forgetting coincide with stable and fast convergence.
@@
Figure 10 studies the influence of batch size on forgetting dynamics. Smaller minibatches lead to higher update variability, resulting in larger oscillations and a greater propensity to forget as training progresses. As batch size increases, efficiency increases and forgetting decreases, plateauing at around one-quarter of the training dataset size. Beyond this, forgetting remains non-zero and is very similar across batch sizes greater than one-quarter of the training dataset.
@@
Finally, Figure 11 examines how model architecture influences forgetting on a high-dimensional classification task (CIFAR-10). We compare logistic regression, a multilayer perceptron (MLP), a convolutional neural network (CNN), and a ResNet. CNNs and ResNets exhibit considerably higher forgetting than logistic regression and MLPs, yet they achieve better task performance. This indicates that the most effective deep learning models do not necessarily minimise forgetting; instead, they maintain a balance where moderate forgetting supports continued adaptation and improved generalisation.
@@
Together, these results (Figure 7, Figure 8, Figure 9, and Figure 10) show that forgetting is a pervasive property of deep learning systems, shaped by both the update dynamics and the learner’s hyperparameters. Consistent with Figure 4, regimes with improved efficiency do not necessarily minimise forgetting. While high levels of forgetting can destabilise training, moderate levels appear beneficial: enabling adaptability while preserving some past information. The relationship between forgetting and learning efficiency is thus nuanced: efficient learning requires not the absence of forgetting but its regulation.
E.2 Forgetting in reinforcement learning
Here, we discuss the impact of hyperparameters in RL on forgetting dynamics.
Buffer size.
Replay buffer size determines which past experience is retained and sampled from during Q-learning updates. A small buffer restricts the effective training distribution to the most recent transitions, while a large buffer maintains long-range temporal support.
Figure 12: Replay buffers regulate prediction support. Forgetting landscapes for different replay buffer sizes. Small buffers produce high, unstable forgetting because the learner continually overwrites supported predictions as the data distribution shifts. Larger buffers stabilise the distribution and reduce forgetting, but overly large buffers reintroduce outdated transitions that also break self-consistency, leading to moderate forgetting and reduced learning efficiency.
With very small buffers, the training distribution shifts rapidly as old transitions are discarded. The learner, therefore, repeatedly overwrites previously supported predictions, producing large and unstable forgetting dynamics. As the buffer size increases, forgetting becomes smaller and more stable: predictions are supported by a more stable training distribution, so self-consistency violations decrease. However, extremely large buffers also reintroduce unsupported transitions, causing the learner to train on outdated targets, resulting in inconsistency. Thus, both forgetting and task performance are optimised at intermediate buffer sizes. This is observed in Figure 12.
Target network update rate.
$\tau$ controls the target Q-network update rate. Therefore, $\tau$ controls how quickly the support for predictions evolves.
Figure 13: Target updates trade-off stability and adaptability. Forgetting landscapes for different target network update rates $\tau$. Large $\tau$ yields less stable forgetting dynamics due to rapidly shifting targets; small $\tau$ produces smooth but uninformative updates that prevent learning. Intermediate $\tau$ values best maintain support for predictions and thus yield balanced forgetting and effective learning.
When $\tau$ is too large, the target network is updated frequently, giving the learner a less stable reference. Predictions can become unsupported more frequently, potentially producing higher, less stable forgetting dynamics. When $\tau$ is very small, the target network is rarely updated. These result in fixed stable targets; however, they also impair the learner’s ability to learn. The learner may not be forgetting, but they are also failing to learn. When $\tau$ is moderate, the target network adjusts gradually, allowing predictions at time $t$ to remain approximately supported by subsequent updates, which is important for self-consistency.
Consequently, both forgetting and performance show optimal behaviour at intermediate $\tau$: the target evolves slowly enough to provide support for predictions, yet quickly enough to adapt the fixed point toward a self-consistent solution (see Figure 13).
Training frequency.
Training frequency determines how often the learner applies updates relative to environment interactions. This determines the rate at which the training distribution shifts.
When the learner makes frequent updates (low training frequency values), the effective training distribution shifts very frequently. As such, updates are frequently performed on shifting distributions, leading to inconsistent predictive distributions over time.
When the learner makes moderately frequent updates, the effective training distribution is quite stable because the learner accumulates small batches of experience before updating. This creates smaller and more stable forgetting as predictions have more persistent support: the effective training distribution changes gradually, and model updates are less likely to be inconsistent.
However, when the learner makes infrequent updates, the effective training distribution again shifts frequently: the learner is updated so infrequently that each update incorporates a highly non-stationary batch of transitions. This again causes model updates to be inconsistent, again breaking support for many of the predictions the learner made earlier, and the forgetting dynamics again increase and become more unstable. In this setting, learning performance also drops because the update frequency impacts the learner’s ability to learn.
Thus, optimal self-consistency emerges only at moderate training frequencies (see Figure 14).
Summary.
Across our RL experiments, we observe that forgetting dynamics are strongly influenced by hyperparameters such as buffer size, target network update rate, $\tau$, and training frequency. Smaller buffer sizes lead to high and unstable forgetting due to insufficient replay support; very low $\tau$ values produce stable forgetting dynamics but prevent meaningful learning; and both low and high training frequencies induce chaotic forgetting dynamics. In all cases, these dynamics previously supported by the learner’s state are no longer reinforced when there is insufficient support, causing higher forgetting.
We also note that forgetting tends to track the TD loss over training, indicating that temporal difference updates implicitly regulate forgetting dynamics. These results highlight that managing forgetting is not just a consequence of network design or training heuristics, but a vital consideration in the design of RL algorithms: poor control of forgetting leads to unstable learning, lower sample efficiency, and suboptimal performance. Optimal algorithm design, therefore, requires balancing the reinforcement of past knowledge with the acquisition of new information.
Figure 14: Training frequency determines forgetting. Forgetting landscapes for different training frequencies. Updating too frequently causes highly unstable forgetting due to rapid shifts in the effective training distribution. Updating too infrequently also causes large shifts in the effective training distribution. A trade-off must be found to achieve more self-consistent updates.